The function f(x;y), defined on pairs of real numbers, satisfies the conditions f(a;a) = 0, f(a; f(b;c)) = f(a;b)+c for any a, b, c. Find f(3.5 ; 7).
1
Expert's answer
2020-12-14T10:31:05-0500

Let "f(a,a)=0" be the condition (1) and "f(a,f(b,c))=f(a,b)+c" be the condition (2).

We have

"f(a,b)+b=f(a,f(b,b))=f(a,0)=f(a,f(a,a))=f(a,a)+a=a", that is "f(a,b)=a-b"

So we obtain that if "f" satisfies the conditions (1) and (2), then "f(a,b)" can be only "a-b"

Check whether "f(a,b)=a-b" satisfies the conditions (1) and (2).

We have "f(a,a)=a-a=0", so "f(a,b)=a-b" satisfies the condition (1).

Next, "f(a,f(b,c))=f(a,b-c)=a-(b-c)=(a-b)+c=f(a,b)+c", so "f(a,b)=a-b" satisfies the condition (2).


We obtain that "f" satisfies the conditions (1) and (2) if and only if "f(a,b)=a-b", then "f(3.5;7)=3.5-7=-3.5"

Answer: "-3.5"